Dorothy is 3 times as old as her sister. In 5 years she will be twice as old as her sister. How old are Dorothy and her sister now?
Answer by OnlineTutor(1)

Let the symbol D stand for Dorothy's current age and the symbol S stand for her sisters current age.
Dorothy is currently 3 times her sisters age. This can be written as:
D = 3 x S
In 5 years Dorothy's age will be D + 5 and her sisters age will be S + 5. At this time Dorothy is twice as old as her sister. Therefore:
(D + 5) = 2 x (S + 5)
D + 5 = 2S + 10
Substitute the first equation into the equation above and you get:
3S + 5 = 2S + 10
S = 5
Substitute S = 5 into the first equation and you get D = 15. That means that Dorothy's current age is 15 and her sisters current age is 5.